a.) 2y′′′−7y′′+12y′+8y=0

b.) y"-2y'+5y=e^xsinx

c.) y′' + y = 2xsinx


1
Expert's answer
2021-11-22T19:27:19-0500

a)


"2y'''-7y''+12y'+8y=0"

Auxiliary equation (or characteristic equation)


"2r^3-7r^2+12r+8=0"

"2r^2(r+\dfrac{1}{2})-8r(r+\dfrac{1}{2})+16(r+\dfrac{1}{2})=0"

"2(r+\dfrac{1}{2})(r^2-4r+8)=0"

"2(r+\dfrac{1}{2})((r-2)^2+4)=0"

"r_1=-\dfrac{1}{2}, r_2=2-2i, r_3=2+2i"

The general solution of the homogeneous differential equation is


"y=c_1e^{-x/2}+e^{2x}(c_2\cos(2x)+c_3\sin(2x))"

b)


"y''-2y'+5y=e^x\sin x"

The corresponding homogeneous differential equation is


"y''-2y'+5y=0"

Auxiliary equation (or characteristic equation)


"r^2-2r+5=0"

"(r-1)^2=-4"

"r_1=1-2i, r_2=1+2i"

The general solution of the homogeneous differential equation is


"y_h=e^{x}(c_1\cos(2x)+c_2\sin(2x))"

Find the particular solution of the nonhomogeneous differential equation in form


"y_p=Ae^x\cos x+Be^x\sin x"

"y_p'=Ae^x\cos x-Ae^x\sin x+Be^x\sin x+Be^x\cos x"

"y_p''=Ae^x\cos x-2Ae^x\sin x-Ae^x\cos x"

"+Be^x\sin x+2Be^x\cos x-Be^x\sin x"

Substitute


"-2Ae^x\sin x+2Be^x\cos x-2Ae^x\cos x"

"+2Ae^x\sin x-2Be^x\sin x-2Be^x\cos x"

"+5Ae^x\cos x+5Be^x\sin x=e^x\sin x"

"3B=1"

"A=0"


"y_p=\dfrac{1}{3}e^x\sin x"

The general solution of the nonhomogeneous differential equation is


"y=e^{x}(c_1\cos(2x)+c_2\sin(2x))+\dfrac{1}{3}e^x\sin x"

c)


"y'' + y = 2x\sin x"

The corresponding homogeneous differential equation is


"y''+y=0"

Auxiliary equation (or characteristic equation)


"r^2+1=0"

"r_1=i, r_2=i"

The general solution of the homogeneous differential equation is


"y_h=c_1\cos x+c_2\sin x"

Find the particular solution of the nonhomogeneous differential equation in form


"y_p=(Ax^2+Bx+C)\cos x+(Dx^2+Ex+F)\sin x"

"y_p'=-(Ax^2+Bx+C)\sin x+(2Ax+B)\cos x"

"+(Dx^2+Ex+F)\cos x+(2Dx+E)\sin x"

"y_p''=-(Ax^2+Bx+C)\cos x-2(2Ax+B)\sin x"

"+2A\cos x-(Dx^2+Ex+F)\sin x"

"+2(2Dx+E)\cos x+2D\sin x"

Substitute


"-(Ax^2+Bx+C)\cos x-2(2Ax+B)\sin x"

"+2A\cos x-(Dx^2+Ex+F)\sin x"

"+2(2Dx+E)\cos x+2D\sin x"

"+(Ax^2+Bx+C)\cos x+(Dx^2+Ex+F)\sin x"

"=2x\sin x"

"-2(2Ax+B)\sin x+2A\cos x"

"+2(2Dx+E)\cos x+2D\sin x=2x\sin x"

"A=-\dfrac{1}{2}"

"B=D=0"

"E=-A=\dfrac{1}{2}"

"y_p=-\dfrac{1}{2}x^2\cos x+\dfrac{1}{2}x\sin x"

The general solution of the nonhomogeneous differential equation is


"y=c_1\cos x+c_2\sin x-\dfrac{1}{2}x^2\cos x+\dfrac{1}{2}x\sin x"